# Canvil: Designerly Adaptation for LLM-Powered User Experiences

### v1

@@ 3. Formative Study
Working with AI, especially LLMs, is an emerging practice in the field of UX (Yang et al., 2020; Yildirim et al., 2023; Feng et al., 2023b; Subramonyam et al., 2021b). We wanted to understand designers’ experiences working on LLM-powered products and features amidst industry-wide shifts towards LLMs (Parnin et al., 2023) as an initial step in our research. We hence conducted a formative study where we interviewed designers about their awareness of, involvement in, and desiderata around designing LLM-powered user experiences.
@@ 3.1. Method and Participants
We conducted 30-minute 1:1 virtual interviews with 12 designers at a large international technology company, where LLM-powered features and products are actively explored. All interviews were conducted in June and July 2023. At the time of the study, all participants were working on products or services that leveraged LLMs in some capacity; LLM application areas spanned conversational search, domain-specific data question-answering (QA), recommendation, text editing and generation, and creativity support tools. Participant details can be found in Table 1 of Appendix A.
@@
  • Desiderata: what do designers desire when crafting LLM-powered user experiences, with regards to both processes and tools?
@@
The first author performed an inductive qualitative analysis of the interview data. This process started with an open coding round in which initial codes were generated, followed by two subsequent rounds of axial coding in which codes were synthesized and merged into higher-level themes. The codes and themes were discussed with research team members at weekly meetings. Other research team members also offered supporting and contrasting perspectives on the codes by writing their own analytic memos, which were also discussed as a team.
@@ 3.2. Findings
We present summaries of our major themes as findings below. While our research did not focus specifically on model adaptation to start, we saw adaptation emerge as a connecting thread across many of our participants’ experiences; as such, we present our findings through the lens of adaptation. Additionally, we use our findings to motivate and define designerly adaptation, culminating in a concrete characterization of the term in Section 3.3.
@@ 3.2.1. Adaptation was seen as a central materialistic property of LLMs
Designers recognized that products delivering compelling, robust user experiences were not powered by out-of-the-box “base” LLMs, but required adaptation and, in some cases, grounding in other sources. P5 gave an example where an out-of-the-box model inappropriately prompted the user for a riddle in a workplace setting: “[The experience] is not quite right, cause it’s in [workplace software] and it’s like, tell me a riddle. I’m at work!” Those working in the domain of conversational search knew the model had to be connected to a search engine to ensure information relevance and reliability, so that they are “not just giving [the user] the chatbot” [P9].
The ability to steer an LLM with natural language enabled designers to envision more flexible and adaptable AI-powered user experiences than working with traditional AI models. Some were familiar with how lightweight adaptation techniques, such as writing system prompts, can be used to specify model behaviors based on desirable UX. P12 discussed an example of how this may work in an entertainment system setting:
“You might want to generate enthusiasm more, right? So the LLM that goes over there might have a UX layer that feels different. It should. I hope it does. The system prompt can say, remember, you’re a machine that wants to get users enthused.” [P12]
@@
Participants recognized that the adaptability of LLMs allows their team to more easily create multiple versions of LLM-powered features in one product. P2 shared that they were already seeing multiple versions of the same model customizing experiences within their product: “There are actually 3 incarnations of the same model, each with slightly different parameters […] based on the prompt that users give, we select which of those 3 incarnations we wanna use.” Participants also believed adapting the model’s ability or behaviors for different user groups or contexts can help create more equitable experiences. For example, P4 and P7 both stated a design goal for their products (in conversational search and domain-specific question-answering, respectively) is to accommodate non-technical users, which can be achieved if the model can “tailor the language depending on technical ability, and gradually introduce [users] to more complicated concepts and terminology.” [P7]. Adaptation thus not only allows user experiences to more flexibly accommodate existing user groups, but also to expand to serving new ones.
3.2.2. Designers were often not involved with adaptation, but desired more involvement.
@@
Even though designers found adaptation inextricably linked to UX considerations, they were typically not involved in adaptation. Some were unsure of how or by whom adaptation is performed and suspected that it was engineers: “I would say [adaptation]’s probably something that was decided by the engineering team […] prompt engineering is more so on the [technical] side of things.” [P2]. Many knew for a fact that it was engineers and data scientists who adapted the models and relied on them for information about model behavior. This often got cumbersome quickly. P11 explained that in their product, designers set character limits for LLM-generated summaries in the UI, but were unable to directly specify and experiment with the model’s output length. Instead, P11 relied on engineers to adapt on their behalf: “I was asking what the difference in summary looks like with 50, 100, and 150 characters. Then [the engineers] would go back and test it and then they would just share the results, like here’s what 50 characters looks like. Here’s what 100 looks like.” P12 experienced a similar procedure and stated that the feedback loop can take as long as a couple weeks. Ideally, they wanted the experimentation to happen “in real time.”
The combination of 1) adaptation and UX being closely intertwined, and 2) designers’ exclusion from the adaptation process, led to a sentiment echoed by the far majority of our participants: designers should have an active role in model adaptation. Indeed, designers are uniquely positioned to contribute to adaptation through their user-centered lens. For example, P4 shared that they naturally gravitate to persona-based reasoning in their design process: “How would I use this? How would my mom use this? Taking on [the perspective of] different personas made sense to me.” This type of reasoning is critical for adapting user experiences to accommodate diverse user needs and usage contexts. Indeed, P10 caught their model’s inability to properly recognize some acronyms users of their product might come across and devised a solution with their team: “I found that [the model] doesn’t know what to do with the acronyms so we floated the idea of having glossary of industry jargon and acronyms [in the system prompt].”
@@
Because of this, many participants were advocates for designers performing adaptation, and considered UX goals as indispensable in this process. P4 shares: “Who better to involve in this process than people whose job it is to think deeply about [the UX]? At the very least, system prompts shouldn’t be written without an understanding of what the end UX goals are.” P9 agreed, saying that designers can offer strategic contributions with user-centered thinking: “It’s super important for designers to think through ideas and make it intuitive for users and help come up with compelling [user] scenarios […] I think design has a bigger opportunity to have a seat at the table strategy-wise.” On a higher level, P12 pointed out that model adaptation can be a contemporary extension of efforts around crafting product voice and tone, which designers are already familiar with: “UX designers and content designers are very attuned to and have pretty much owned the story around voice and tone of products, and have for years and years.”
We see a broad consensus among participants that there is both demand and opportunity for designers to engage in adaptation. We therefore propose that designerly adaptation should leverage designers’ existing expertise and workflows to enhance model behavior in user-centered ways.
3.2.3. Designers lack support to tinker with and adapt LLMs
@@
Overall, participants felt limited by current tools and resources to tinker with LLMs. These findings echo prior work that finds that the inability to directly access the models and experiment with their capabilities and limitations (i.e., “tinkering”) is a primary challenge in the design process for AI-powered user experiences (Yang et al., 2020; Subramonyam et al., 2022). Many participants emphasized the importance of “tinkering” to understand model capabilities and limitations to inform design ideation and interface prototyping. After some hands-on experience with the model, P2 said they could much more clearly “see or gauge the power of the language model,” while P9 tinkered with a model they had early access to and recalled that “you can identify some gaps [in capability] right off the bat.”
While some designers welcomed the easy access to ChatGPT or tried out the GPT playground as a means to familiarize themselves with the “base” LLMs, participants recognized the necessity of tinkering with the adapted models. To gain access, designers either had to keep burdening the engineers or wait until a test version of the product is launched. P3 shared that “The only way we would have [to tinker] is using the [product] online and play around with what [the team] did.” P10 expressed frustration at this workflow: “You want to be able to play with [the model] yourself and understand what the user experience is like. And I can’t play with it.
Some designers took the initiative to seek out new tools to tinker with and adapt models—currently, these tools are mostly limited to either programming with APIs or playground interfaces with many technical developer settings. However, usage of those tools did not persist. P1 tried various tools within and outside of their company and did not find them to be designer-friendly: “[The tools are] still kind of technical, a lot of [designers] don’t realize how parameters work. Like how does temperature work?” Others, such as P4, did not find the tools well-suited for their use case: “I haven’t seen anything that feels like it makes it really easy to make small changes to a [system] prompt and then run the same set of queries over it and see whether it makes a difference.”
Taken together, we envision adaptation and tinkering must be tightly coupled for designers to work with adapted LLMs within their design process. This includes being able to easily create, test, and iterate on versions of adaptation, and having this workflow be integrated into their design environment to reduce friction and eliminate dependence on engineers. Moreover, as noted in Section 2.2, some approaches to adaptation demand technical skills designers typically do not acquire in their training, and also do not fit within the tight feedback loops of the design process due to high data and compute overheads. Thus, we consider prompt-based adaptation methods to be most ideal for designerly adaptation. This broadens participation for designers while still allowing for modification of many aspects of model behavior.
3.3. Summary: A Characterization of Designerly Adaptation
We present three characteristics of designerly adaptation through several key insights drawn from our formative study’s findings. First, not all forms of adaptation can be considered as designerly adaptation. Designerly adaptation should not demand a high degree of technical expertise in AI—instead, designers should be capable of and prepared to conduct it with their current skillset and training. One promising approach to designerly adaptation, then, is to author system prompts with natural language. Second, designerly adaptation should take advantage of designers’ unique expertise and perspectives to inform model behavior. This may include incorporating user research into the system prompts. Third, tinkering should have a major role in designerly adaptation, encouraging tight feedback loops of creation, testing, and iteration that align with feedback loops in the broader design process.
4. Canvil: a Technology Probe for Designerly Adaptation
In light of our formative study, we introduce Canvil, [^5] The name “Canvil” is a portmanteau of “canvas” and “anvil.” We envision Canvil to be a metaphorical anvil by which LLM behavior can be shaped within design canvases. a technology probe for designerly adaptation in the form of a Figma widget. We first outline our design goals, which we ground in our formative study’s findings. We then walk through Canvil’s user interface and implementation.
@@ 4.1. Design Goals
Our design goals, informed primarily by insights from our formative study and supported by our literature review, guided why and how we built Canvil. They are as follows.
@@
    Support designerly adaptation via system prompting.
    Our formative study revealed that some designers were already experimenting with system prompting in their design workflows (Sections 3.2.1 and 3.2.2). System prompts are also an appropriate medium for architecting UX—they shape a model’s behaviors across all or a specified group of user inputs. We also note that among the diverse techniques for adaptation (see Section 2.2), system prompting is conducted in natural language, making it accessible to designers, many of whom lack a technical background.
@@
    Leverage abstractions and environments designers are already familiar with.
    Designers who want to be more involved in adaptation can feel unsupported due to existing tools (e.g., calling APIs programmatically) not aligning with their established workflows and mental models. Indeed, we saw in Section 3.2.3 that designers sought out new tools for adaptation but discarded them after a while because they were too technical or did not match desired use cases. Given that designers prefer to stay in one tool of their choice (e.g., Figma) for most stages of the design process (Feng et al., 2023c), we hypothesize that tools for designerly adaptation should be tightly integrated with design tools and leverage common abstractions within those tools (e.g., layers, components, frames) to better align with designers’ mental models.
@@
    Allow for user research to inform model behavior.
    When an LLM is integrated into a user-facing application, it becomes part of a sociotechnical system embedded in the deployment context (Selbst et al., 2019). Designers are well-positioned to understand this context and leverage their understanding for LLM adaptation from working closely with users or user research data (Section 3.2.2). Therefore, we invite designers to consider the user and social factors surrounding the UX at hand in tandem with the technical capabilities of the model. That is, we empower designers to adapt model behavior with user research insights.
@@
    UX is a highly collaborative practice (Feng et al., 2023c; Deng et al., 2023). In our formative study, we saw that designers collaborated closely with not only other designers, but also product managers, software engineers, data scientists, and more. All these roles can contribute to adaptation. Moreover, not all desired model behaviors can be realized with prompt-based adaptation alone—for example, connecting an LLM to a knowledge base will likely require collaboration with a developer or data scientist (Section 3.2.1). We therefore design our system to afford collaboration with diverse technical and non-technical stakeholders.
  • DG5:
    Enable tightly coupled model adaptation and tinkering within the design process.
    Designers found model tinkering highly valuable in our formative study as it allowed them to better understand LLMs as a design material—how they behave, what they are capable of, and where their limitations are. This aligned with prior work on designers working with AI models more generally (Subramonyam et al., 2021a; Feng and McDonald, 2023). Yet, our formative study indicates that designers currently had limited to no involvement in adaptation, nor access to tinkering with adapted models. Even if some limited opportunities for tinkering exist, they would require designers to leave their established workflows to seize those opportunities. We thus aim to incorporate tinkering and adaptation seamlessly into the design process.
@@ 4.2. Canvil as a Probe
We envision designerly adaptation as a new opportunity for designers to contribute to model adaptation efforts. To investigate whether and how this new practice may be included in existing design workflows, we designed Canvil as a technology probe.
Commonly used in contextual research in HCI (Jörke et al., 2023; Hohman et al., 2019), a technology probe is an artifact, typically in the form of a functional prototype (Hutchinson et al., 2003), presented to the user “not to capture what is so much as to inspire what might be” (Lury and Wakeford, 2012). That is, probes offer one instantiation of tooling and/or interaction techniques for a domain to better understand phenomena within that domain. Hutchinson et al. (Hutchinson et al., 2003) state that technology probes have three goals: the social science goal of understanding users in a real-world context, the engineering goal of field-testing the technology, and the design goal of inspiring new technologies. We map these three goals onto our objectives with Canvil:
@@
  • Engineering: develop and launch a Figma widget that connects to OpenAI’s LLMs and allows them to be adapted via system prompting from within Figma.

### v2

@@ 3. Formative Study
Working with AI, especially LLMs, is an emerging practice in the field of UX (Yang et al., 2020; Yildirim et al., 2023; Feng et al., 2023b; Subramonyam et al., 2021b). We conducted a formative interview study to understand designers’ experiences working on LLM-powered products and features amidst industry-wide shifts towards LLMs in 2023 (Parnin et al., 2023) as an initial step in our research.
@@ 3.1. Method and Participants
We conducted 30-minute 1:1 virtual interviews with 12 designers at a large international technology company where LLM-powered features and products are actively explored. All interviews were conducted in June and July 2023. At the time of the study, all participants were working on products or services that leveraged LLMs in some capacity; LLM application areas spanned conversational search, question-answering (QA) on domain-specific data, recommendation, text editing and generation, and creativity support tools. Participant details can be found in Table 1 of Appendix A.
@@
  • Desiderata: what do designers desire when crafting LLM-powered UX, with regards to both processes and tools?
@@
The first author performed an inductive qualitative analysis of the interview data. This process started with an open coding round in which initial codes were generated, followed by two subsequent rounds of axial coding in which codes were synthesized and merged into higher-level themes. The codes and themes were discussed with research team members at weekly meetings. Other research team members also offered supporting and contrasting perspectives on the codes by writing their own analytical memos, which were also discussed as a team.
@@ 3.2.1. Adaptation was seen as a central materialistic property of LLMs
Participants recognized that products delivering compelling, robust UX were not powered by out-of-the-box “base” LLMs, but required adaptation based on user needs and contexts. P5 gave an example where an out-of-the-box model inappropriately prompted the user for a riddle in a workplace setting: “[The experience] is not quite right, cause it’s in [workplace software] and it’s like, tell me a riddle. I’m at work!” P4 and P7 both stated a design goal for their products (in conversational search and domain-specific question-answering, respectively) is to accommodate non-technical users, which can be achieved if the model can “tailor the language depending on technical ability, and gradually introduce [users] to more complicated concepts and terminology” [P7]. P2 shared that their team settled on using three adapted versions of the same model to customize UX within their product: “There are actually three incarnations of the same model, each with slightly different parameters […] based on the prompt that users give, we select which of those 3 incarnations we wanna use.”
In their design work, participants appreciated the ability to steer an LLM with natural language to envision more flexible and adaptable AI-powered UX. Some were familiar with how lightweight adaptation techniques, such as writing system prompts, can be used to specify model behaviors based on design specifications. P12 discussed an example of how this may work in an entertainment system setting:
“You might want to generate enthusiasm more, right? So the LLM that goes over there might have a UX layer that feels different. The system prompt can say, remember, you’re a machine that wants to get users enthused.” [P12]
P12’s approach points to an important distinction in outcome evaluation for when a model is adapted for design purposes versus conventional prompt engineering: the end solution to a design problem is not a prompt, but a user experience. In P12’s case, they opted for the model to act with enthusiasm in their product. Whether the model can generate enthusiastic responses and whether the model delivers a desirable user experience within the product are two different (albeit potentially related) questions. The latter likely also depends on design choices made elsewhere in the product will be evaluated through user studies rather than model performance metrics.
3.2.2. Designers envisioned constraints and desiderata for model behavior, but were unable to directly apply them
@@
Participants often proposed constraints and desiderata for model behavior in their design work. For example, P11 explained that in their product, designers set character limits for LLM-generated summaries in the UI, while P5 and P12 both created “personalities” for the LLM in their products. However, participants rarely had the opportunity to tinker with and apply their desiderata and constrints themselves, resulting in overly cumbersome workflows. P11 shared that they relied on engineers to adapt on their behalf: “I was asking what the difference in summary looks like with 50, 100, and 150 characters. Then [the engineers] would go back and test it and then they would just share the results, like here’s what 50 characters looks like. Here’s what 100 looks like.” P12 experienced a similar procedure and stated that the feedback loop can take as long as a couple weeks. Ideally, they wanted the experimentation to happen “in real time.” In general, designers were typically not involved in model adaptation decisions. Some were unsure of how or by whom adaptation is performed and suspected that it was engineers: “I would say [adaptation]’s probably something that was decided by the engineering team […] prompt engineering is more so on the [technical] side of things.” [P2].
@@
Because model behavior and UX are inextricably linked, many participants desired more involvement in steering models directly, particularly towards UX goals. P4 shares: “Who better to involve in this process than people whose job it is to think deeply about [the UX]? At the very least, system prompts shouldn’t be written without an understanding of what the end UX goals are.” P9 agreed, saying that designers can offer strategic contributions with user-centered thinking: “It’s super important for designers to think through ideas and make it intuitive for users and help come up with compelling [user] scenarios […] I think design has a bigger opportunity to have a seat at the table strategy-wise.” On a higher level, P12 pointed out that model adaptation can be a contemporary extension of efforts around crafting product voice and tone, which designers are already familiar with: “UX designers and content designers are very attuned to and have pretty much owned the story around voice and tone of products, and have for years and years.”
We see strong evidence showing that designers are uniquely positioned to contribute to adaptation through their user-centered lens. Although they may primarily adapt models via prompt-based methods, their contributions extend far beyond prompt engineering. First, they define a key prerequisite for prompt engineering: the UX goals which the model should be steered towards and evaluated against. This is not a one-way process—these goals may be modified based on hands-on experimentation with the model. However, we did not observe a two-way process in practice as designers were often not involved in adaptation themselves. Second, they can surface additional constraints and desiderata for prompt engineered models through UX evaluation and testing. Indeed, P10 caught their model’s inability to properly recognize some acronyms in their product during testing and devised a solution with their team: “I found that [the model] doesn’t know what to do with the acronyms so we floated the idea of having glossary of industry jargon and acronyms [in the system prompt].” In sum, designers’ user-centered insights help improve LLM-powered UX, addressing crucial decisions in both goal-setting before prompt engineering and UX evaluation after the fact.
3.2.3. Designers did not find sufficient support in current LLM tools and resources
@@
Overall, participants felt limited by current tools and resources for tinkering with LLMs within the design process. Prior work found that the inability to directly access the models and experiment with their capabilities and limitations is a primary challenge in the design process for AI-powered UX (Yang et al., 2020; Subramonyam et al., 2022). This was reflected in our findings as well—in the few cases where designers had hands-on access to model tinkering, designers recalled the experience being highly valuable. P2 said they could much more clearly “see or gauge the power of the language model,” while P9 tinkered with a model they had early access to and recalled that “you can identify some gaps [in capability] right off the bat.” However, most did not have access to tinkering and found this frustrating. Designers either had to keep burdening the engineers or wait until a test version of the product is launched. P3 shared that “The only way we would have [to tinker] is using the [product] online and play around with what [the team] did.” Similarly, P10 felt limited by this workflow: “You want to be able to play with [the model] yourself and understand what the UX is like. And I can’t play with it.
Some designers took the initiative to seek out new tools to tinker with LLMs. Designers welcomed the easy access to free LLM chatbots such as ChatGPT, but did not feel like they were well-integrated into their design workflows. Many tried a variety of tools—mostly playground interfaces for prompt engineering—without much success. Among others, P1 did not find these interfaces to be designer-friendly: “[The tools are] still kind of technical, a lot of [designers] don’t realize how many parameters work. Like how does temperature work?” Others wanted these tools to leverage the same metaphors as already-familiar design tools, such as “a UX library with some specific patterns you can use” [P4] or “accessing content right away in a [design system]” [P6].
Our findings reveal fundamental differences in tooling needs between designers who adapt LLMs for UX design purposes and prompt engineers. For designers, the model is a design material through which they explore design solutions. That is, unlike prompt engineering, the material itself is not the final product—the UX is. The UX does not stop nor end with model behavior. For example, designers will also need to consider user journeys leading up to the LLM interaction, and subsequent user interactions with the LLM output as well as other UI components Designers work with many different types of materials within the broader design process, and typically have adequate tools to understand and manipulate these materials. For example, designers can easily draft wireframes and create animations in modern design tools such as Figma. When it comes to understanding and manipulating LLMs, designers are much less well-supported. Prompt engineering tools do not satisfy designers’ needs because they help craft prompts rather than UX and they do not fit into existing design processes.
3.3. Summary: The Case For Designerly Adaptation
In his 1982 essay, Nigel Cross characterized the uniqueness of epistemological inquiry in the field of design by outlining several aspects of designerly way of knowing (Cross, 1982). One key aspect was the use of design codes. To Cross, design codes offer a two-way translation between abstract design requirements and concrete designed objects, where the objects are a form of knowledge about how to satisfy certain requirements (Cross, 1982)—that is, codes represent the entanglement of a design process with its design materials. The process generates requirements that, when applied to the materials, results in a designed object, while the properties of the material may simultaneously shape the process used to work with it. For some designers in our formative study, LLM prompts provided an accessible means of expressing design requirements in natural language to adapt model behavior. However, prompting alone did not offer the entanglement needed to shape a designed object (Section 3.2.2), nor did it allow LLMs’ material properties to influence designers’ practices due to prompting tools’ lack of integration with design work (Section 3.2.3). Although LLMs present a new and exciting material for designers, there was no design code to entangle the material with their existing processes.
We thus propose a code, which we call designerly adaptation. We define designerly adaptation to be a translational mechanism by which designers entangle process with material by translating requirements into designs for LLM-powered UX—and vice versa—through adapting LLM behavior. Our proposed code is designerly in nature because the resulting object (the LLM-powered UX) embodies design knowledge constructed from intersecting processes with materials. To further characterize designerly adaptation, we present four design goals to guide its practical implementation.
@@ 3.4. Design Goals
Prompting does not serve as a satisfactory design code because 1) it is often divorced from design practices, environments, and abstractions; and 2) it is unclear how prompts can be used to express design requirements. These realizations informed our design goals below. Our goals guide our efforts to operationalize designerly adaptation with a design tool.
@@
    Integrate seamlessly with existing design environments and abstractions.
    Designerly adaptation is, at its core, a design practice. Given designers’ preference to stay in one tool of their choice (e.g., Figma) for most stages of the design process (Feng et al., 2023c), we aim for Canvil to smoothly integrate with existing design tools and leverage common abstractions within those tools (e.g., layers, components, frames) to align with designers’ mental models. Indeed, we saw designers discarding tools like ChatGPT and prompt playgrounds because they did not fit into their design workflows or were overly technical (3.2.3).
@@
    Support direct translation of design requirements into model behavior.
    As noted in Sections 3.2.2 and 3.3, designers lacked the codes to directly translate and apply envisioned constraints in their design work. This undermines designers’ contributions to LLM-powered products—after all, they are well-positioned to understand this context using user-centered methods and synthesize their understanding and findings into constraints on model behavior (3.2.2). With designerly adaptation and Canvil, we provide a new avenue of translation between requirements and LLM-powered UX to empower designers’ participation in model adaptation.
@@
    Enable iterative model tinkering within the design process.
    Designers found model tinkering highly valuable in our formative study as it allowed them to quickly understand LLMs as a design material—how they behave, what they are capable of, and where their limitations are (3.2.1), aligning with findings from prior work (Subramonyam et al., 2021a; Feng and McDonald, 2023). A design code that encourages tinkering can therefore support rapid ideation, testing, and iteration of many possible solutions and foster productive co-evolution between design requirements and designed objects.
@@
    UX practice is highly collaborative (Feng et al., 2023c; Deng et al., 2023). Designers told us that they collaborated closely with not only other designers, but also product managers, software engineers, data scientists, and more (3.2.1). Designerly adaptation may be primarily performed by designers but should welcome collaboration with non-design stakeholders. This not only allows more diverse insights to be incorporated into adaptation, but also catalyzes further refinement of adapted models (particularly by technical teams) towards production readiness.
4. Canvil: a Technology Probe for Designerly Adaptation
In light of our formative study and design goals, we introduce Canvil, [^4] The name “Canvil” is a portmanteau of “canvas” and “anvil.” We envision Canvil to be a metaphorical anvil by which LLM behavior can be shaped within design canvases. a technology probe for designerly adaptation in the form of a Figma widget. In this section, we describe Canvil’s design choices, user interface, and implementation.
@@ 4.1. Canvil as a Probe
We designed Canvil as a technology probe to investigate whether and how designerly adaptation may be included in existing design workflows. Commonly used in contextual research in HCI (Jörke et al., 2023; Hohman et al., 2019), a technology probe is an artifact, typically in the form of a functional prototype (Hutchinson et al., 2003), presented to the user “not to capture what is so much as to inspire what might be” (Lury and Wakeford, 2012). That is, probes offer one instantiation of tooling and/or interaction techniques for a domain to better understand phenomena within that domain. Hutchinson et al. (Hutchinson et al., 2003) state that technology probes have three goals: the social science goal of understanding users in a real-world context, the engineering goal of field-testing the technology, and the design goal of inspiring new technologies. We map these three goals onto our objectives with Canvil:
@@
  • Engineering: develop and launch a Figma widget that connects to LLMs via APIs and allows them to be adapted and used from within Figma.